### v4

# Designing quantum chemistry algorithms with just-in-time compilation

Abstract.
We introduce just-in-time (JIT) compilation to the integral kernels for Gaussian-type orbitals (GTOs) to enhance the efficiency of electron repulsion integral computations. For Coulomb and exchange (JK) matrices, JIT-based algorithms yield a 2× speedup for the small 6-31G* basis set over GPU4PySCF v1.4 on an NVIDIA A100-80G GPU. By incorporating a novel algorithm designed for orbitals with high angular momentum, the efficiency of JK evaluations with the large def2-TZVPP basis set is improved by up to 4×. The core CUDA implementation is compact, comprising only ~1,000 lines of code, including support for single-precision arithmetic. Furthermore, the single-precision implementation achieves a 3× speedup over the previous state-of-the-art.
1. Introduction
Traditional quantum chemistry software (LSZC, 25; WSP+, 25; WHJ+, 24) has historically relied on ahead-of-time (AOT) compilation and monolithic code generation, with tightly coupled Fortran or C++ implementations (Nee, 22; EGF+, 21; FTS+, 16) tuned for a fixed set of use cases. Although this design has served well for decades, it poses serious limitations in some modern high-performance computing (HPC) environments where GPUs are the primary source of computational power. Algorithms in electronic structure theory often require handling a combinatorial variety of tensor contraction patterns and sparsity conditions that depend on the input molecular system. Static compilation must conservatively support all conditional branches, hyperparameter combinations, and numerical thresholds, leading to bloated binaries, excessive control flow divergence, poor cache and register utilization, and inefficient use of memory bandwidth. On GPUs, these inefficiencies are further magnified by strict constraints on register pressure, warp divergence, and hardware occupancy, often requiring significant manual optimization. For example, most quantum chemistry packages deal with low-angular-momentum integrals successfully but still struggle with the performance of high-angular momentum integrals, which suffer from all of these problems.
Just-in-time (JIT) compilation offers a transformative solution by enabling runtime code specialization based on the actual inputs: e.g., integral angular-momentum pattern, the number of primitives in an orbital shell, and task type. By generating code on demand, JIT techniques can eliminate irrelevant branches, tailor memory access patterns, and fuse computational stages that would be separate in statically compiled workflows. In GPU architectures, this dynamic specialization translates into reduced register pressure, improved memory coalescing, and better occupancy. Similar ideas have seen widespread adoption in the machine learning community, where frameworks such as PyTorch’s TorchScript (PGM+, 19), Triton (TKC, 19) and JAX (BFH+, 18) leverage JIT to fuse operations, reduce memory movement, and generate hardware-specific kernels. These successes demonstrate the maturity and impact of JIT compilation as a general-purpose performance strategy in high-throughput numerical computing, and motivate its application to quantum chemistry workloads.
Beyond runtime performance, just-in-time compilation also offers substantial advantages to developers. JIT facilitates rapid prototyping and testing: since code is generated at runtime, one can easily benchmark different kernel variants, tiling strategies, or instruction sequences, without recompiling and linking the entire application. This accelerates development cycles, simplifies debugging, and lowers the barrier to contributing performance improvements across architectures. Furthermore, the developers are able to design novel algorithms which are not general enough under the AOT framework. In this paper, we demonstrate several algorithms due to this benefit. The development experience can be even further improved: Within a narrow area of application such as machine learning, SQL and GPU programming, one can design domain-specific languages (DSLs) via LLVM (LA, 04). Frameworks like TensorFlow(AAB+, 15), PyTorch(PGM+, 19), JAX(BFH+, 18), and TVM incorporate their own DSLs—such as TensorFlow’s GraphDef and XLA HLO, PyTorch’s TorchScript and FX, JAX’s JAXpr, and TVM’s TIR—to represent algorithms in intermediate forms that enable aggressive optimization and efficient execution. These DSLs allow developers to write high-level code while benefiting from low-level performance tuning through JIT compilation and backend-specific code generation, bridging the gap between productivity and performance.
JIT compilation benefits mixed precision algorithms in quantum chemistry, which naturally accelerate electron repulsion integrals by exploiting the Schwarz inequality (LUM, 11; MM, 13). Most tasks can be calculated in single-precision (FP32) as compared to the commonly-used double-precision (FP64). Single precision algorithms benefit quantum-chemistry computations in three key ways: (1) deep learning inference GPUs and consumer-grade GPUs typically feature far more FP32 units than FP64 units, consume less power, and are more readily available; (2) single-precision data occupy half the storage, allowing more values to reside in shared memory and registers, improving data locality and throughput; and (3) hardware-accelerated FP32 implementations of the exponential (exp) and error function (erf) offer 4×–8× speedups, alleviating what is a computational bottleneck in certain workloads. These benefits can be better utilized by JIT compilation without significant changes to the code.
Contributions: 1) We introduce JIT to the quantum chemistry community and optimize the two-electron integral kernels (1q1t in Sec. 3.2) with compilation techniques. Our framework integrates seamlessly with the quantum chemistry package GPU4PySCF (LSZC, 25; WSP+, 25) and is released as the open-source xQC library. 2) We develop a novel fragmentation algorithm for high angular momentum integrals, improving data locality and alleviating memory-bandwidth bottlenecks through multilevel reduction. 3) We benchmark single-precision integral kernels under our JIT framework and demonstrate substantial performance gains when combining JIT compilation with our new algorithmic strategies.
2. A short formulation of Rys quadrature
Evaluation of the electron repulsion integral is one major bottleneck of density functional theory (DFT), the most widely used electronic structure theory (Bec, 14). In this paper, the two-electron repulsion integrals are calculated with the Rys quadrature algorithm (FL, 08). Other popular integral algorithms such as McMurchie–Davidson (MD, 78) and Head-Gordon-Pople (HP, 88) have been implemented in (AV, 24; WHJ+, 24; VBa, 23). Reimplementation of these methods with JIT is out of the scope of this paper. In this section, we briefly formulate the Rys quadrature algorithm for clarifying the notations and data dependency used in the rest of the paper. For detailed derivation and numerical stability analysis, we refer the readers to (FL, 08).
Here we are using notations instead of conventional notations to index atomic orbitals, to be consistent with the pseudo-code later in this paper. Since each of the Gaussian type orbitals is contracted from a set of primitive Gaussian functions (HJO, 00), the integral tensor for a shell quartet is expressed as the sum of the primitive integral tensors
, , , and are the pre-defined contraction coefficients for each shell in the basis set. The number of Gaussian functions in each primitive shell is or for cartesian atomic orbitals or spherical atomic orbitals respectively, where is the angular momentum of the shell. Only Cartesian orbitals are considered in the following sections, and spherical orbitals are just a linear combination of Cartesian orbitals with fixed transformation coefficients. These linear transformations can be performed efficiently outside the kernel. As a result, each dimension of the primitive integral is . We denote the dimension of the integral for each shell quartet as .
Within Rys quadrature formalism, each primitive integral can be evaluated as a weighted sum of the integrant at specific quadrature root points
where is the number of Rys roots. and are the roots and weights of Rys quadrature. The roots and weights depend on the exponents and center (atomic) positions of the four primitive Gaussian functions, the total angular momentum, and a range-separation parameter if needed. The calculation of Rys roots and weights relies on interpolations from tabulated data, and the details have been documented in (FL, 08). The number of Rys roots is determined at the compilation time, and only the data(WSP+, 25; Sun, 15) associated with the number of roots are compiled into the binary. The primitive integral at each root position can be factored into intermediate variables in three spatial directions:
| for any | ||||
, , and are the intermediate variables introduced in (DRK, 76; FL, 08). Each of them is a tensor of shape . They depend on the exponents, atomic coordinates and angular momentum of the four primitive Gaussians only. They can be efficiently evaluated using recursive relations, documented in (FL, 08). But storing those values on the limited resources on GPU can be tricky, and several strategies will be discussed in Sec. 3.1 and Sec. 3.2. The index set for enumerating the Guassian functions of a primitive shell is defined as
The idx function is defined differently in various quantum chemistry packages. One will need to double check the conventions to align the quantities in atomic basis, such as density matrix, molecular orbitals, and so on. We adopt the convention of reversed lexicographic order. The indexing operation in general makes the memory access random and less efficient.
Because of its huge size, the four-dimensional integral tensor is not formed explicitly in the main storage, but instead contracted with the density matrix to form the Coulomb () and exchange () potential matrices:
For a molecular system with real-valued orbitals, the 8-fold symmetry of can be used to save computational cost. In particular, for the integral element composed of the -th, -th, -th and -th atomic orbital, it is contracted with two density matrix elements and to form two matrix elements and , and it is contracted with four density matrix elements , . , to form four matrix elements , , , .
Conventionally, the Schwarz inequality for shell quartets
is used to screen the significant shell quartets, i.e. all integrals within an insignificant shell quartet are considered negligible and not evaluated. The cost of generating the significant shell quartet is almost negligible, even for the low-angular-momentum integrals. Other screening techniques can be combined with the algorithms. They will be discussed in section 6.
In the direct self-consistent field (SCF) algorithm, the integrals are further screened by the density matrix
and
where , , and are the changes of density matrix, Coulomb potential, and exchange potential between two steps of SCF iterations (i.e. incremental SCF), respectively. Since is getting smaller along the SCF iterations, the screening is more effective. In practice, the density matrix screening is performed at shell pair level, which is similar to Eq. (5).
3. Design and Algorithms
The input of a quantum chemistry calculation contains the protocol of the method, geometry of a molecule, and various thresholds. These inputs are unknown when a typical quantum chemistry package is compiled. Therefore, all types of input have to be handled as dynamic variables. The idea of JIT design is to feed the input information to the compiler at compilation time such that the compiler can aggressively optimize the code. On the other hand, if the compilation completely depends on all the input information, i.e. if we recompile for every job even if successively jobs are only different in geometry, the recompilation may slow down the overall execution time, since the compilation can be more expensive than the actual calculation with an improper design. In the design of JIT programs, distinguishing between static and dynamic variables is crucial for user experience and performance. In general, we follow this principle: The information of quantum chemistry protocol and device information are static for compilation. That is, when the quantum chemistry method, basis set, or type of GPUs is changed, the kernels should be recompiled. The same kernels will be used for different molecules. All the molecule information, including the atomic numbers, atomic types and atomic coordinates, and the total charge and spin of the system, are dynamic variables for compilation.
More specifically, we select the angular momentum and contraction pattern of the basis set, as well as the task type, as the template parameters for the kernels. They are described in Table 1. Once these variables are fixed, the size of any intermediate variable in the integral kernels is known at compilation time. If those parameters are changed, we have to recompile the kernels. With such choices, we avoid complicated runtime logic flows for different data type, task type, and unrolling strategies. Thus the amount of source code is minimal: only ~1,000 lines of code, much less than the ~20,000 lines in GPU4PySCF v1.4.
Data Type | Description | ||
| int | Angular momentum of shells | |||
| do_j | bool | Whether to compute the J matrix | ||
| do_k | bool | Whether to compute the K matrix | ||
| rys_type | int | 0: no range separation (full range), | ||
| 1: long-range, -1: short-range | ||||
| n_dm | int | # of density matrices | ||
| (e.g. two for unrestricted DFT) | ||||
| DataType | type | float or double precision | ||
| int | # of primitives in the | |||
| (i, j, k, l) shells | ||||
| int | Fragment sizes of the | |||
| shells (details in Sec. 3.2) |
The parameters of the basis set are dynamic variables. Since many elements share the same contraction pattern but different coefficients and exponents, we do not need to compile multiple kernels for the same contraction pattern. Generally, reading those data is not the bottleneck of the kernels. All of the dynamic parameters are summarized in 2.
Data Type | Description | |
| double/single | contraction coefficients in basis | ||
| double/single | exponents in basis | ||
| coords | double/single | coordinates of atoms | |
| double | density matrix as input | ||
| double | coulomb potential as output | ||
| double | exchange potential as output | ||
| int | # of atomic orbitals | ||
| double | range-separation parameter |
An example of the DFT self-consistent field (SCF) workflow is presented in Fig. 1. The workflow employs NVRTC (NVI, 25), the runtime compilation library for CUDA, via CuPy(OUN+, 17) as JIT compiler. NVRTC offers the direct control over shared memory, as well as a relatively fast compilation. Other compilers, such as Numba, can be employed similarly with minimal modification, which will be discussed in Sec. 6
3.1. One quartet one thread (1q1t) algorithm
The first algorithm we present is largely the same as the classical Rys quadrature algorithm on CPUs (Sun, 15), except for a more aggressive loop unrolling strategy is used. First, a list of significant shell quartets is generated by running over all possible combinations screened by the Schwartz inequality. Each quartet is assigned to one thread for evaluating all elements of the integral in Eq. (1) (including all primitives in a contraction and all functions in a shell), and contract it with the density matrix according to Eq. (4), until the pool of all significant shell quartet is empty. This algorithm was implemented in the early version of GPU4PySCF (v1.0 and before). This straightforward implementation has two performance issues. 1) There are still multiple loops not unrolled due to unknown variables at compilation time, such as number of primitivies for each shell and number of density matrices. This is actually the main bottleneck of low-angular-momentum integrals. 2) Significant register spill out onto local memory for high angular momentum. In this subsection, we will show that with more known variables at compilation time, the performance is significantly improved. The second issue will be discussed in the next subsection.
With the design in this work, when handling one shell quartet, the ranges of all loops are known at compilation time. The JIT compiler can aggressively unroll those loops, especially loops for primitives, and use registers more efficiently.
In the single-precision implementation, the exponents, contraction coefficients, and atomic coordinates are prepared and loaded in FP32. The density matrix values are loaded as FP64 and then casted to FP32 in kernels. Inside the kernel, all the arithmetic operations are performed in FP32 and accelerated with the --use_fast_math compiler flag. One should notice that fast-math intrinsics for single precision is much faster, but not IEEE-754 accurate. When and are calculated, they are first casted to FP64, then accumulated with atomicAdd in FP64.
We benchmark the performance of JK kernels with a straight peptide chain composed of 120 glycines (gly120), and 6-31G (DP, 75) basis set. Density matrix screening is turned off for profiling purpose. The timings on A100-80G and A10-24G are showned in Figure 2 and 3 respectively. The overall speedup of FP64 against AOH compilation in GPU4PySCF v1.4 is about 2.0x. For low-angular-momentum, the evaluations of exp and erf functions are often the computational bottleneck. Thanks to the special hardware acceleration of these functions on GPU, the overall speedup of FP32 reaches about 5-10x on A100-80G. Since A10-24G has 1/32 hardware units for FP64 compared to FP32, FP64 algorithms are always bounded by the limited compute units. The improvement with JIT is marginal. When switched to FP32, the 1q1t algorithm is accelerated by 30x.
There are still several corner-case kernels. When the number of primitives is greater than 5, the NVRTC compiler refuses to unroll the nest loops, even if we specify #pragma unroll. Performance drop is observed. The issue will be discussed in Appendix (Section B). Another case is in double precision which requires 96 registers per thread for the intermediate variables, and 162 registers for the integral elements. The kernel consumes at least 258 registers, which already exceeds on the hard limit per thread (255 registers) by the CUDA platform. The register spill affects the performance dramatically. Higher angular momentum integrals suffer from the same issues. They will be handled by different algorithm later in Section 3.2.
3.2. Quartet fragmentation (1qnt) algorithm
For high-angular momentum integrals, the intermediate variables and integral elements of one shell quartet cannot fit in the registers. From GPU4PySCF v1.0-v1.4, a fixed-size 1D fragmentation strategy is introduced to split the integral into multiple threads. Instead of one block thread processing all integrals of one shell quartet, a group of 2 to 256 threads cooperatively compute all integral elements within one shell quartet. The strategy significantly improved the performance of the high-angular integrals. In this work, we introduce the dynamic fragmentation strategy to further improve the data locality, and multi-level reduction algorithm to reduce the bottleneck of memory access (both shared memory and global memory). Since multiple threads for the same shell quartet use the same intermediate variables, the intermediate variable , and will be stored in shared memory. Then, each thread in the group will compute and store into registers only a fraction of the integrals. For example, the shape of the integral (dd—dd) is (6,6,6,6), each thread computes a fragment with the size of (3,3,3,3). 16 threads are needed to store the entire (dd—dd) integral. In the reduction stage, each thread is responsible for a fragment of 3x3 and . The fragments and are further reduced at the thread block level. The reduction scheme is presented in Algorithm 1. The shape of the intermediate variables are also known at compilation time. A similar idea has been proposed in (AAF+, 10), where is partitioned in , , and indices. is partitioned in and indices.
Optimize the fragmentation scheme. The performance of the integral kernel is sensitive to the fragment size. Different fragmentation strategies can easily lead to a 10x difference in performance. The fragment size can impact the register usage of each thread, the shared memory usage of each thread block, and the number of shell quartet assigned to each block. The resource consumption also need to meet the hard requirements of CUDA compute capability. We have to take these factors into account for searching the optimal fragment size. This optimization problem is less likely to be solved analytically. We look for the optimal fragmentation strategy by grid search. The searching space of high angular momentum can be huge. We made several assumptions:
We assume that each dimension of the integral is a multiple of the fragment size. There probably exists a better tiling strategy without this constraint. However, based on the grid search results, we found that the divisible fragment size, in general, is optimal. An additional benefit of this constraint is that each thread does not need to check the bound of iteration, which saves the significant computational cost.
The optimal fragment size is independent of the number of primitives. This assumption in general is not true, since the kernel with more number of primitives take more registers. Then, less registers are left for the integral fragment. The assumption is made for compromising with the high computational cost. We search the optimal fragment size based on a fixed number of primitives for low-angular orbital (s, p orbitals). The number of primitives for high-angular atomic orbital (d, f, g orbitals) is set to be 1.
The block size is always 256 threads.
Since the double-precision and single-precision algorithms differ in their usage of shared memory, register count, and hardware units, a strategy optimized for double precision is not necessarily optimal for single precision. Therefore, we optimize fragmentation strategies separately for each precision. The default fragmentation schemes are tuned on NVIDIA A100-80G and A10-24G GPUs, respectively. To ensure compatibility, the fragment size is optimized under the constraint of 48 KB shared memory per streaming multiprocessor, allowing the strategy to be applicable to older GPU generations. An optimization script is included in the xQC release, enabling users to retune the fragment size for different GPU architectures if they diverge significantly from the default targets.
We benchmark the 1tnq algorithm with a straight peptide chain composed of 30 glycines (gly30) and def2-TZVPP (WA, 05) basis set. Density matrix screening is turned off. Results on NVIDIA A100-80G and A10-24G GPUs appear in Figure 4 and 5, respectively. Using the default fragmentation for FP64, the 1tnq algorithm is 5-8x faster than GPU4PySCF v1.4 on A100-80G. When switched to FP32, the algorithm is further accelerated by 2x. For high-angular-momentum integrals, the computational cost of exp and erf functions is not significant relative to multiply/add operations. On A10-24G, the high-angular-momentum integrals are accelerated by roughly 2x for FP64. This is due to the algorithm is GPU4PySCF v1.4 is memory bound. The 1qnt algorithm reduces the memory bottleneck by multi-level reduction. When switched to FP32, the integrals are accelerated by 10-20x. Further performance analysis is provided in Appendix (Section B).
4. Results
4.1. Compilation Overhead
Just-in-time compilation introduces non-negligible overhead, but this cost is amortized through on-disk caching and in-memory reuse. When a kernel is first invoked, it is compiled and the resulting binary is saved to disk. The cached binaries can be reused across the process. During the first SCF iteration, each process loads these binaries into RAM directly; all subsequent iterations then fetch them directly from memory. Table 3 summarizes the one-time compilation times for various kernels using CUDA v12.1. As expected, high–angular-momentum integrals take longer to compile and generate larger binaries: for the C, H, O, and N elements with the 6-31G* basis (140 kernels), the compilation time is just 30 s and the total kernel size is roughly 10 MB, while the def2-TZVPP basis (750 kernels) yields about 100 MB of binary and takes approximately 200 s to compile. Once those kernels are compiled, subsequent calls take less than 1 s to load cached binaries.
These figures serve as a reference. Actual compilation times will vary with different CUDA versions and hardware quality, particularly disk performance. On modern devices, the compilation overhead can be reduced in various ways. In practice, especially in homogeneous cloud environments, it can be advantageous to store compiled kernels on shared or remote storage and download them as needed. The similar approach has been used in the machine learning community, e.g. PyTorch AOTInductor. Alternatively, retaining intermediate .ptx files and loading them directly allows users to bypass one step of compilation.
| Integral pattern | Algorithm |
| (ss—ss)/(11—11) | 1Q1T | 284.8 | 25 | ||||
| (ss—ss)/(33—33) | 1Q1T | 512.0 | 74 | ||||
| (pp—pp)/(11—11) | 1Q1T | 638.8 | 75 | ||||
| (pp—pp)/(33—33) | 1Q1T | 880.6 | 94 | ||||
| (dd—dd)/(11—11) | 1QnT | 6011 | 194 | ||||
| (ff—ff)/(11—11) | 1QnT | 7697 | 270 | ||||
| (gg—gg)/(11—11) | 1QnT | 59468 | 602 |
4.2. Benchmark with GPU4PySCF and TeraChem
In this work, integrals for the Coulomb (J) and exchange (K) terms are screened simultaneously. GPU4PySCF and xQC use the same screening technique. The number of significant shell quartets in each JK calculation is identical in both cases. In double-precision, xQC reproduces GPU4PySCF’s results exactly, so we omit a detailed FP64 accuracy analysis. Under our single precision scheme, JK matrix elements are computed in single precision but accumulated in FP64; all other operations including diagonalization and direct inversion of the iterative subspace (DIIS)—remain in FP64.
We use a customized TeraChem v1.9 build with f-orbital support. Although both GPU4PySCF and TeraChem employ the Schwarz inequality to bound integral magnitudes, they apply it at different levels: TeraChem at the primitive shell level and GPU4PySCF at the contracted level. Their SCF initial guesses and DIIS settings also differ, so our GPU4PySCF–TeraChem comparisons are not strictly apples-to-apples. For all calculations shown, the screening threshold for both J and K matrices is set to . SCF with GPU4PySCF/xQC converges in 14 iterations, whereas the same SCF with TeraChem converges in 11 iterations. This does not imply TeraChem’s settings are superior, as convergence behavior can vary with the system. A sample TeraChem input is provided in Appendix A. All the basis sets are in Cartesian format.
When both codes run in FP64, their total energies agree to within 0.001 mHa. When all integrals are evaluated in FP32, we observe uncertainties on the order of 1 mHa in both xQC and TeraChem (Table 4), reinforcing the well-known result that pure single-precision arithmetic is insufficient for high-accuracy quantum-chemical calculations. A more practical mixed-precision algorithm is among our future directions.
| Molecule | Basis |
| gly30 | 6-31G* | 0.23 mHa | 0.0005 mHa | -0.07 mHa | ||||||
| gly30 | def2-TZVPP | 1.95 mHa | 0.0004 mHa | 2.10 mHa | ||||||
| valinomycin | 6-31G* | -0.06 mHa | -0.001 mHa | -0.26 mHa | ||||||
| valinomycin | def2-TZVPP | 0.68 mHa | -0.001 mHa | 0.63 mHa |
As shown in Table 5, xQC delivers a twofold speedup over GPU4PySCF v1.4 when using the 6-31G* basis set and a fourfold gain with the larger def2-TZVPP basis. Although TeraChem v1.9 generally outpaces GPU4PySCF v1.4, its advantage is only marginal. In double-precision, xQC outperforms TeraChem by about 4x when def2-TZVPP basis set is used, and a less significant 1.6x for 6-31G* basis. In single-precision, xQC slightly outperforms TeraChem on the small basis set and runs twice as fast on def2-TZVPP. Both TeraChem and xQC still employ double precision for eigenvalue decomposition and DIIS, making these linear-algebra routines a significant cost on the A10-24 GB GPU. Isolating the JK kernel, xQC is approximately three times faster than TeraChem.
| Molecule | Protocol | GPU4PySCF | TeraChem | xQC | |||
| JK/cycle | SCF | JK/cycle | SCF | JK/cycle | SCF | ||
| gly30 | HF/6-31G*/A100-80G/FP64 | 1.9 s | 28.1 s | 1.3 s | 19.08 s | 0.8 s | 12.6 s |
| gly30 | HF/def2-TZVPP/A100-80G/FP64 | 78.4 s | 1125.7 s | 91.4 s | 1055.1 s | 18.8 s | 268.1 s |
| gly30 | HF/6-31G*/A10-24G/FP32 | - | - | 1.0 s | 23.12 s | 0.7 s | 18.08 s |
| gly30 | HF/def2-TZVPP/A10-24G/FP32 | - | - | 63.5 s | 870.81 s | 21.2 s | 423.19 s |
5. Conclusion
We optimized the low–angular-momentum electron repulsion integral algorithm in GPU4PySCF v1.4 using just-in-time (JIT) compilation and introduced a novel high–angular-momentum integral method. The core CUDA implementation for both algorithms spans only 1,000 lines of code in total. JIT compilation not only accelerates the evaluation of Gaussian-type-orbital integrals but also provides developers with greater flexibility. Our streamlined implementation delivers a 2-4x performance improvement over GPU4PySCF v1.4 and is significantly easier to maintain. Additionally, it supports single- and double-precision arithmetic, accepts multiple density matrices as input, and integrates seamlessly with the existing GPU4PySCF framework.
6. Future Work and Limitations
High level abstractions of JIT. Although our current implementation relies heavily on handwritten CUDA kernels, it would be preferable to express these routines in a high-level language such as Triton (TKC, 19), Numba (LPS, 15), or JAX (BFH+, 18). These frameworks offer richer support for kernel fusion, making it easier to embed quantum-chemistry operations within machine-learning pipelines. For example, one could fuse the on-the-fly computation of integrals with subsequent matrix–matrix multiplications (not only contraction with density matrices), yielding more efficient end-to-end workflows. Further, one could write a customized kernel for the derivatives with respect to contraction coefficients and exponents, then port to an existing auto-differentiation framework (ZC, 22).
Supporting more functionalities. A practical quantum‐chemistry workflow requires more than the acceleration of Coulomb and exchange kernels. The exchange-correlation functional in DFT, the gradients and Hessians of Coulomb, exchange and exchange-correlation terms, post Hartree-Fock methods, and a wide array of molecular properties are all essential. By integrating xQC with GPU4PySCF, users can build sophisticated simulation pipelines, and any of these routines can be further accelerated on demand using JIT compilation.
Mixed precision algorithms. This paper focuses on the performance of the double precision and single precision separately. Thanks to the Schwartz inequality, the computational load of the integrals can be naturally split into FP64 calculations and FP32 calculations without loss of accuracy theoretically. Other research efforts (LUM, 11; MM, 13) show that 90% of the computations can be in FP32. A naive mixed precision strategy can be easily proposed based on this work. One should notice, even though the value of integral can be represented in FP32, not every underlying arithmetic operation for computing the integral can necessarily be computed with FP32. More effort is needed to rigorously implement a strategy without error accumulation.
Screening techniques. The intrinsic scalings of the Coulomb and exchange terms are different. The number of significant Coulomb terms increases quadratically, while the number of significant exchange terms increases linearly with respect to the system size. It is more efficient to evaluate the Coulomb and exchange terms separately. In some quantum chemistry packages such as TeraChem, the Coulomb terms are efficiently evaluated with J-engine (WH, 96). Besides the Schwartz inequality, which gives a rigorous upper bound, other estimates can further eliminate the unnecessary calculations. For example, the MBIE (LDO, 05), QQR (MLFO, 12), and CSAM (TO, 17). MBIE provides a rigorous upperbound, but it is expensive to calcuate. QQR and CSAM take advantage of distance dependence in their estimates, hence improving the performance for large systems. But those estimates are less rigorous.
Author Contributions
Xiaojie Wu led the conceptualization of JIT algorithms and developed xQC v0.1, and wrote the initial draft. Qiming Sun contributed the improvements of the JK kernels v1.0 to v1.4 in GPU4PySCF, which are not documented elsewhere. Yuanheng Wang explored single-precision algorithms in GPU4PySCF, discussed JIT algorithms in xQC v0.1, and wrote the manuscript.
Appendix A TeraChem input sample
In this paper, we use the following sample TeraChem script. The coordinates, basis, and precision arguments are changed correspondingly.
run energy method hf basis 6-31gs precision single coordinates gly30.xyz charge 0 timings yes threcl 1e-13 threex 1e-13 guess sad xtol 1e-6 purify no fock incremental
Appendix B Analyze the bottlenecks of 1q1t algorithm and 1qnt algorithm
We benchmark the 1q1t algorithm on the gly120 system with the following customized basis set in NWChem format (PAD+, 19),
X S
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
0.2700058226E+00 1
The same basis set is assigned to all types of elements, for profiling purpose only. The number of primitives varies in the tests. The 1q1t algorithm is largely memory bound when the number of primitives is small. For a (ss—ss) kernel with uncontracted orbitals, when switching from FP64 to FP32, the speedup of FP64/FP32 is only 2. The computing intensity increases rapidly with the number of primitives. When the numbers of primitives of all the shells are 4, the FP32 to FP64 speedup achieves the theoretical peak performance of A10-24G, which has 32:1 FP64 / FP32 compute units. However, when the number of primitives is greater than 5, we observed the register spill out onto local memory. This is essentially due to the compiler (NVRTC/NVCC) not unrolling the nested loops of primitives. There are multiple ways to manually unroll the loops, but for simplicity of the code, we leave it for future work. The similar issue also occurs in the case of other patterns, such as , , and so on.
The performance bottleneck in the 1qnt algorithm arises from two distinct phases: integral generation and contraction. In the integral generation phase, the cost of calculating the integral values is independent of the number of density matrices. In the contraction phase, the integral value is contracted with every density matrix, so its cost grows linearly with the matrix count. To quantify this behavior, we vary the number of density matrices and recorded the total wall-clock time (Figure 7). We constructed the following basis set in NWChem format for the test, which assigns the same d and f orbtials to all types of elements:
X D
1.2700058226E+00 1
X F
1.2700058226E+00 1
The results show an almost perfectly linear relation between the overall cost and the number of density matrices. For the (dd—dd) pattern, integral generation requires 96 ms and each contraction 67 ms; in a higher-angular-momentum case, these times rise to 1,518 ms and 837 ms, respectively. Consequently, when only one density matrix is used (e.g., in restricted HF/DFT methods), the bottleneck lies in integral generation, whereas with two or more matrices (for example, unrestricted HF/DFT methods), density matrix contraction dominates. In the latter scenario, more sophisticated contraction strategies are needed, such as tensor core-accelerated kernels.
We emphasize that this analysis is approximate: increased register pressure from handling multiple density matrices can further affect kernel performance, and those effects are not reflected in our current timing estimates.
Acknowledgements.
The authors appreciate the support of Dr. Wen Yan and the ByteDance Seed management team. Special thanks are also extended to Dr. Zhichen Pu for his discussions. The authors also thank the PySCF community for providing valuable tools that contributed to this work, especially the GPU4PySCF contributors.References
- AAB+ [15] Martín Abadi, Ashish Agarwal, Paul Barham, Eugene Brevdo, Zhifeng Chen, Craig Citro, Greg S. Corrado, Andy Davis, Jeffrey Dean, Matthieu Devin, Sanjay Ghemawat, Ian Goodfellow, Andrew Harp, Geoffrey Irving, Michael Isard, Yangqing Jia, Rafal Jozefowicz, Lukasz Kaiser, Manjunath Kudlur, Josh Levenberg, Dandelion Mané, Rajat Monga, Sherry Moore, Derek Murray, Chris Olah, Mike Schuster, Jonathon Shlens, Benoit Steiner, Ilya Sutskever, Kunal Talwar, Paul Tucker, Vincent Vanhoucke, Vijay Vasudevan, Fernanda Viégas, Oriol Vinyals, Pete Warden, Martin Wattenberg, Martin Wicke, Yuan Yu, and Xiaoqiang Zheng. TensorFlow: Large-scale machine learning on heterogeneous systems, 2015. Software available from tensorflow.org.
- AAF+ [10] Andrey Asadchev, Veerendra Allada, Jacob Felder, Brett M. Bode, Mark S. Gordon, and Theresa L. Windus. Uncontracted rys quadrature implementation of up to g functions on graphical processing units. Journal of Chemical Theory and Computation, 6(3):696–704, 2010. PMID: 26613300.
- AV [24] Andrey Asadchev and Edward F. Valeev. 3-center and 4-center 2-particle gaussian ao integrals on modern accelerated processors. The Journal of Chemical Physics, 160(24):244109, 06 2024.
- Bec [14] Axel D. Becke. Perspective: Fifty years of density-functional theory in chemical physics. The Journal of Chemical Physics, 140(18):18A301, 04 2014.
- BFH+ [18] James Bradbury, Roy Frostig, Peter Hawkins, Matthew James Johnson, Chris Leary, Dougal Maclaurin, George Necula, Adam Paszke, Jake VanderPlas, Skye Wanderman-Milne, and Qiao Zhang. JAX: composable transformations of Python+NumPy programs, 2018.
- DP [75] James D. Dill and John A. Pople. Self-consistent molecular orbital methods. xv. extended gaussian-type basis sets for lithium, beryllium, and boron. J. Chem. Phys., 62:2921–2923, 1975.
- DRK [76] Michel Dupuis, John Rys, and Harry F. King. Evaluation of molecular integrals over gaussian basis functions. The Journal of Chemical Physics, 65(1):111–116, 07 1976.
- EGF+ [21] Evgeny Epifanovsky, Andrew T. B. Gilbert, Xintian Feng, Joonho Lee, Yuezhi Mao, Narbe Mardirossian, Pavel Pokhilko, Alec F. White, Marc P. Coons, Adrian L. Dempwolff, Zhengting Gan, Diptarka Hait, Paul R. Horn, Leif D. Jacobson, Ilya Kaliman, Jörg Kussmann, Adrian W. Lange, Ka Un Lao, Daniel S. Levine, Jie Liu, Simon C. McKenzie, Adrian F. Morrison, Kaushik D. Nanda, Felix Plasser, Dirk R. Rehn, Marta L. Vidal, Zhi-Qiang You, Ying Zhu, Bushra Alam, Benjamin J. Albrecht, Abdulrahman Aldossary, Ethan Alguire, Josefine H. Andersen, Vishikh Athavale, Dennis Barton, Khadiza Begam, Andrew Behn, Nicole Bellonzi, Yves A. Bernard, Eric J. Berquist, Hugh G. A. Burton, Abel Carreras, Kevin Carter-Fenk, Romit Chakraborty, Alan D. Chien, Kristina D. Closser, Vale Cofer-Shabica, Saswata Dasgupta, Marc de Wergifosse, Jia Deng, Michael Diedenhofen, Hainam Do, Sebastian Ehlert, Po-Tung Fang, Shervin Fatehi, Qingguo Feng, Triet Friedhoff, James Gayvert, Qinghui Ge, Gergely Gidofalvi, Matthew Goldey, Joe Gomes, Cristina E. González-Espinoza, Sahil Gulania, Anastasia O. Gunina, Magnus W. D. Hanson-Heine, Phillip H. P. Harbach, Andreas Hauser, Michael F. Herbst, Mario Hernández Vera, Manuel Hodecker, Zachary C. Holden, Shannon Houck, Xunkun Huang, Kerwin Hui, Bang C. Huynh, Maxim Ivanov, Ádám Jász, Hyunjun Ji, Hanjie Jiang, Benjamin Kaduk, Sven Kähler, Kirill Khistyaev, Jaehoon Kim, Gergely Kis, Phil Klunzinger, Zsuzsanna Koczor-Benda, Joong Hoon Koh, Dimitri Kosenkov, Laura Koulias, Tim Kowalczyk, Caroline M. Krauter, Karl Kue, Alexander Kunitsa, Thomas Kus, István Ladjánszki, Arie Landau, Keith V. Lawler, Daniel Lefrancois, Susi Lehtola, Run R. Li, Yi-Pei Li, Jiashu Liang, Marcus Liebenthal, Hung-Hsuan Lin, You-Sheng Lin, Fenglai Liu, Kuan-Yu Liu, Matthias Loipersberger, Arne Luenser, Aaditya Manjanath, Prashant Manohar, Erum Mansoor, Sam F. Manzer, Shan-Ping Mao, Aleksandr V. Marenich, Thomas Markovich, Stephen Mason, Simon A. Maurer, Peter F. McLaughlin, Maximilian F. S. J. Menger, Jan-Michael Mewes, Stefanie A. Mewes, Pierpaolo Morgante, J. Wayne Mullinax, Katherine J. Oosterbaan, Garrette Paran, Alexander C. Paul, Suranjan K. Paul, Fabijan Pavošević, Zheng Pei, Stefan Prager, Emil I. Proynov, Ádám Rák, Eloy Ramos-Cordoba, Bhaskar Rana, Alan E. Rask, Adam Rettig, Ryan M. Richard, Fazle Rob, Elliot Rossomme, Tarek Scheele, Maximilian Scheurer, Matthias Schneider, Nickolai Sergueev, Shaama M. Sharada, Wojciech Skomorowski, David W. Small, Christopher J. Stein, Yu-Chuan Su, Eric J. Sundstrom, Zhen Tao, Jonathan Thirman, Gábor J. Tornai, Takashi Tsuchimochi, Norm M. Tubman, Srimukh Prasad Veccham, Oleg Vydrov, Jan Wenzel, Jon Witte, Atsushi Yamada, Kun Yao, Sina Yeganeh, Shane R. Yost, Alexander Zech, Igor Ying Zhang, Xing Zhang, Yu Zhang, Dmitry Zuev, Alán Aspuru-Guzik, Alexis T. Bell, Nicholas A. Besley, Ksenia B. Bravaya, Bernard R. Brooks, David Casanova, Jeng-Da Chai, Sonia Coriani, Christopher J. Cramer, György Cserey, III DePrince, A. Eugene, Jr. DiStasio, Robert A., Andreas Dreuw, Barry D. Dunietz, Thomas R. Furlani, III Goddard, William A., Sharon Hammes-Schiffer, Teresa Head-Gordon, Warren J. Hehre, Chao-Ping Hsu, Thomas-C. Jagau, Yousung Jung, Andreas Klamt, Jing Kong, Daniel S. Lambrecht, WanZhen Liang, Nicholas J. Mayhall, C. William McCurdy, Jeffrey B. Neaton, Christian Ochsenfeld, John A. Parkhill, Roberto Peverati, Vitaly A. Rassolov, Yihan Shao, Lyudmila V. Slipchenko, Tim Stauch, Ryan P. Steele, Joseph E. Subotnik, Alex J. W. Thom, Alexandre Tkatchenko, Donald G. Truhlar, Troy Van Voorhis, Tomasz A. Wesolowski, K. Birgitta Whaley, III Woodcock, H. Lee, Paul M. Zimmerman, Shirin Faraji, Peter M. W. Gill, Martin Head-Gordon, John M. Herbert, and Anna I. Krylov. Software for the frontiers of quantum chemistry: An overview of developments in the q-chem 5 package. The Journal of Chemical Physics, 155(8):084801, 08 2021.
- FL [08] N. Flocke and V. Lotrich. Efficient electronic integrals and their generalized derivatives for object oriented implementations of electronic structure calculations. Journal of Computational Chemistry, 29(16):2722–2736, 2008.
- FTS+ [16] M. J. Frisch, G. W. Trucks, H. B. Schlegel, G. E. Scuseria, M. A. Robb, J. R. Cheeseman, G. Scalmani, V. Barone, G. A. Petersson, H. Nakatsuji, X. Li, M. Caricato, A. V. Marenich, J. Bloino, B. G. Janesko, R. Gomperts, B. Mennucci, H. P. Hratchian, J. V. Ortiz, A. F. Izmaylov, J. L. Sonnenberg, D. Williams-Young, F. Ding, F. Lipparini, F. Egidi, J. Goings, B. Peng, A. Petrone, T. Henderson, D. Ranasinghe, V. G. Zakrzewski, J. Gao, N. Rega, G. Zheng, W. Liang, M. Hada, M. Ehara, K. Toyota, R. Fukuda, J. Hasegawa, M. Ishida, T. Nakajima, Y. Honda, O. Kitao, H. Nakai, T. Vreven, K. Throssell, J. A. Montgomery, Jr., J. E. Peralta, F. Ogliaro, M. J. Bearpark, J. J. Heyd, E. N. Brothers, K. N. Kudin, V. N. Staroverov, T. A. Keith, R. Kobayashi, J. Normand, K. Raghavachari, A. P. Rendell, J. C. Burant, S. S. Iyengar, J. Tomasi, M. Cossi, J. M. Millam, M. Klene, C. Adamo, R. Cammi, J. W. Ochterski, R. L. Martin, K. Morokuma, O. Farkas, J. B. Foresman, and D. J. Fox. Gaussian˜16 Revision C.01, 2016. Gaussian Inc. Wallingford CT.
- HJO [00] Trygve Helgaker, Poul Jørgensen, and Jeppe Olsen. Atomic Basis Functions, chapter 6, pages 201–255. John Wiley & Sons, Ltd, 2000.
- HP [88] Martin Head‐Gordon and John A. Pople. A method for two‐electron gaussian integral and integral derivative evaluation using recurrence relations. The Journal of Chemical Physics, 89(9):5777–5786, 11 1988.
- LA [04] C. Lattner and V. Adve. Llvm: a compilation framework for lifelong program analysis & transformation. In International Symposium on Code Generation and Optimization, 2004. CGO 2004., pages 75–86, 2004.
- LDO [05] Daniel S. Lambrecht, Bernd Doser, and Christian Ochsenfeld. Rigorous integral screening for electron correlation methods. The Journal of Chemical Physics, 123(18):184102, 11 2005.
- LPS [15] Siu Kwan Lam, Antoine Pitrou, and Stanley Seibert. Numba: a llvm-based python jit compiler. In Proceedings of the Second Workshop on the LLVM Compiler Infrastructure in HPC, LLVM ’15, New York, NY, USA, 2015. Association for Computing Machinery.
- LSZC [25] Rui Li, Qiming Sun, Xing Zhang, and Garnet Kin-Lic Chan. Introducing gpu acceleration into the python-based simulations of chemistry framework. The Journal of Physical Chemistry A, 129(5):1459–1468, 2025. PMID: 39846468.
- LUM [11] Nathan Luehr, Ivan S. Ufimtsev, and Todd J. Martínez. Dynamic precision for electron repulsion integral evaluation on graphical processing units (gpus). Journal of Chemical Theory and Computation, 7(4):949–954, 2011. PMID: 26606344.
- MD [78] Larry E McMurchie and Ernest R Davidson. One- and two-electron integrals over cartesian gaussian functions. Journal of Computational Physics, 26(2):218–231, 1978.
- MLFO [12] Simon A. Maurer, Daniel S. Lambrecht, Denis Flaig, and Christian Ochsenfeld. Distance-dependent schwarz-based integral estimates for two-electron integrals: Reliable tightness vs. rigorous upper bounds. The Journal of Chemical Physics, 136(14):144107, 04 2012.
- MM [13] Yipu Miao and Kenneth M. Jr. Merz. Acceleration of electron repulsion integral evaluation on graphics processing units via use of recurrence relations. Journal of Chemical Theory and Computation, 9(2):965–976, 2013. PMID: 26588740.
- Nee [22] F. Neese. Software update: the orca program system, version 5.0. WIRES Comput. Molec. Sci., 12(1):e1606, 2022.
- NVI [25] NVIDIA Corporation. NVRTC: NVIDIA Runtime Compilation Library. NVIDIA Corporation, 2025. Accessed: 2025-06-17.
- OUN+ [17] Ryosuke Okuta, Yuya Unno, Daisuke Nishino, Shohei Hido, and Crissman Loomis. Cupy: A numpy-compatible library for nvidia gpu calculations. In Proceedings of Workshop on Machine Learning Systems (LearningSys) in The Thirty-first Annual Conference on Neural Information Processing Systems (NIPS), 2017.
- PAD+ [19] Benjamin P. Pritchard, Doaa Altarawy, Brett Didier, Tara D. Gibsom, and Theresa L. Windus. A new basis set exchange: An open, up-to-date resource for the molecular sciences community. J. Chem. Inf. Model., 59:4814–4820, 2019.
- PGM+ [19] Adam Paszke, Sam Gross, Francisco Massa, Adam Lerer, James Bradbury, Gregory Chanan, Trevor Killeen, Zeming Lin, Natalia Gimelshein, Luca Antiga, Alban Desmaison, Andreas Köpf, Edward Yang, Zach DeVito, Martin Raison, Alykhan Tejani, Sasank Chilamkurthy, Benoit Steiner, Lu Fang, Junjie Bai, and Soumith Chintala. PyTorch: an imperative style, high-performance deep learning library. Curran Associates Inc., Red Hook, NY, USA, 2019.
- Sun [15] Qiming Sun. Libcint: An efficient general integral library for gaussian basis functions. Journal of Computational Chemistry, 36(22):1664–1671, 2015.
- TKC [19] Philippe Tillet, H. T. Kung, and David Cox. Triton: an intermediate language and compiler for tiled neural network computations. In Proceedings of the 3rd ACM SIGPLAN International Workshop on Machine Learning and Programming Languages, MAPL 2019, page 10–19, New York, NY, USA, 2019. Association for Computing Machinery.
- TO [17] Travis H. Thompson and Christian Ochsenfeld. Distance-including rigorous upper bounds and tight estimates for two-electron integrals over long- and short-range operators. The Journal of Chemical Physics, 147(14):144101, 10 2017.
- VBa [23] Jorge Luis Galvez Vallejo, Giuseppe M.J. Barca, and Mark S. Gordon and. High-performance gpu-accelerated evaluation of electron repulsion integrals. Molecular Physics, 121(9-10):e2112987, 2023.
- WA [05] Florian Weigend and Reinhart Ahlrichs. Balanced basis sets of split valence, triple zeta valence and quadruple zeta valence quality for h to rn: Design and assessment of accuracy. Phys. Chem. Chem. Phys., 7:3297, 2005.
- WH [96] Christopher A. White and Martin Head‐Gordon. A j matrix engine for density functional theory calculations. The Journal of Chemical Physics, 104(7):2620–2629, 02 1996.
- WHJ+ [24] Yuanheng Wang, Diptarka Hait, K. Grace Johnson, O. Jonathan Fajen, Juncheng Harry Zhang, Rubén D. Guerrero, and Todd J. Martínez. Extending gpu-accelerated gaussian integrals in the terachem software package to f type orbitals: Implementation and applications. The Journal of Chemical Physics, 161(17):174118, 11 2024.
- WSP+ [25] Xiaojie Wu, Qiming Sun, Zhichen Pu, Tianze Zheng, Wenzhi Ma, Wen Yan, Yu Xia, Zhengxiao Wu, Mian Huo, Xiang Li, Weiluo Ren, Sheng Gong, Yumin Zhang, and Weihao Gao. Enhancing gpu-acceleration in the python-based simulations of chemistry frameworks. WIREs Computational Molecular Science, 15(2):e70008, 2025. e70008 CMS-1146.R2.
- ZC [22] Xing Zhang and Garnet Kin-Lic Chan. Differentiable quantum chemistry with pyscf for molecules and materials at the mean-field level and beyond. The Journal of Chemical Physics, 157(20):204801, 11 2022.